# Three-dimensional shape and connectivity of physical networks

### v1

@@ IV.2 Link volume and shape
where $V_{\textnormal{seg}}(v,w)$ is the volume of each segment tracing the link $(i,j)$. We normalize the link volume by the total volume of the network $V_{\textrm{total}}$, setting the unit of measurement. We find that linked trees, or the fruit fly neural networks, have consistently high link volume heterogeneity, as their distributions span 6 to 10 orders of magnitude, which is higher compared to most lattices and trees (Fig.4a).
We also measure the aspect ratio of physical links $a(i,j)=\rho_{\textrm{link}}(i,j)/l_{\textrm{link}}(i,j)$, where $\rho_{\textrm{link}}$ is the average radius and $l_{\textrm{link}}(i,j)$ is the length of link $(i,j)$. We find the largest average aspect ratios for one of the fruit fly networks and the anthill ($med(a)=\frac{1}{3}$), while $med(a)$ is substantially lower for other data sets (See SI Sec. I A). Overall, this confirms that physical links are elongated tube-like objects.
Since physical links are tube-like objects, we can capture most of their shape by characterizing their one-dimensional trajectory. Here, we calculate the deviation of the link trajectories from a straight line, quantifying how curved a link is. For this we rely on a measure of straightness introduced originally in the context of geographical networks [5], namely we calculate the complimentary straightness for each link
@@
Calculating the median of complimentary link straightness distribution, $\text{med}(\bar{S})$, reveals that links in all of the 15 physical networks tend to follow a trajectory close to a straight line: most networks have $\text{med}(\bar{S})\approx 0.1$, meaning that the length of links is most often less than 10% longer than the optimal straight trajectory. Similarly to link volume heterogeneity, linked trees tend to cluster together and are among the networks with the straightest links with $\text{med}(\bar{S})=0.05$. Reference [21] introduced random linear physical networks, a minimal model that constructs physical network from straight cylinders. The fact that we observed an abundance of straight or close-to-straight links lends support for using such linear physical network models to understand the role of physicality in real networks. Note, however, that although most links are close to straight, the distribution of $\bar{S}$ is right-skewed as seen in Fig. 4, which points to a smaller fraction of links that significantly deviate from a straight trajectory. (See SI Sec. III and SI Sec. IV D).
@@
Finally, we computed the correlations between link straightness $S(i,j)$ and the total link length $l_{\textrm{link}}(i,j)$ and volume $V_{\textrm{link}}(i,j)$ for each dataset using Kendall’s rank correlation coefficient $\tau$ [1] Figure 4b shows that for all networks, there is a positive rank correlation $\tau>0$ between $\bar{S}(i,j)$ and $l_{\textrm{link}}(i,j)$, indicating that longer links tend to follow a more winding path. We also observe a positive correlation $\tau>0$ between $\bar{S}(i,j)$ and $V_{\textrm{link}}(i,j)$, since longer links tend to have larger volume. The only exception to this are the fruit fly neural networks, for which $\bar{S}(i,j)$ and $V_{\textrm{link}}(i,j)$ are negatively correlated $\tau<0$. To explain this, note that fruit fly neural networks are composed of neurons that have large somas, which are represented in the dataset as short, yet high-volume segments, hence the largest volume of physical links are short and straight (See SI Sec. I G).
@@
Figure 4: Link shape, length and volume. (a) Box plots of $\bar{S}$ and $V_{\textrm{link}}$, where the vertical lines indicate the median, the boxes span the $25$th-$75$th percentile range, and the whiskers extend from the $0.1$th to the $99.9$th percentile. We find that the bulk of the distributions fall in the range between $\textrm{med}(\bar{S})=0$ and $\textrm{med}(\bar{S})=0.1$, indicating that the networks are mostly composed of close-to-straight links. Link volume $V_{\textrm{link}}$ distributions span a wider range for linked trees, and are the most narrow for lattices. (b) By computing Kendall rank correlation $\tau$ between link complementary straightness $\bar{S}$, link volume $V_{\textrm{link}}$ and link trajectory length $l_{\textrm{link}}$, we observe a consistent trend of $\tau>0$, meaning that longer and more voluminous links tend to have more winding paths. This trend is only reversed for the fruit fly networks, which have $\tau<0$ between $\bar{S}$ and $V_{\textrm{link}}$: neuron somas are represented as high-volume, straight links composed of a small number of skeleton segments.
@@ IV.3 Link confinement
In the previous sections, we found that space-filling and link properties are heterogeneously distributed: most regions of space are sparse and most links are straight, yet there exists dense regions of the network and a small fraction of links follow paths that deviate from a straight line significantly. This suggests that volume exclusion or other repulsive physical interactions may also play an uneven role in shaping the network. To further investigate this hypothesis we devise a quantity that captures the confinement of a link by other components of a network.
@@
The trajectory of a physical link $(i,j)$ in our datasets is given by the ordered set ${\mathcal{T}}(i,j)$ composed of a series of oriented three-dimensional segments. To generate the random trajectory ${\mathcal{T}}_{\textrm{r}}(i,j)$, we shuffle the order of the segments while maintaining their orientation and length, creating a uniform random permutation of ${\mathcal{T}}(i,j)$. The randomization preserves the endpoints and the total length of the link, but otherwise removes any correlation between subsequent segments; therefore, the possible link trajectories ${\mathcal{T}}_{\textrm{r}}(i,j)$ have the same complimentary straightness $\bar{S}$ as the original link. Next, we estimate $I(i,j;l,k)$, the expected number of intersections between the randomized link ${\mathcal{T}}_{\textrm{r}}(i,j)$ and each other non-randomized link ${\mathcal{T}}(k,l)$. To quantify the confinement of the link $(i,j)$, we sum up the expected number of intersections with other links:
@@
where the first term corresponds to intersections when link $(i,j)$ is randomized, and the second term corresponds to intersections when link $(l,k)$ is randomized. Note that the summation in Eq. (6) excludes links that share an endpoint with $(i,j)$. We do this to exclude trivial intersections form the count, since adjacent links $(i,j)$ and $(j,k)$ necessarily overlap at the junction point $j$ even for non-randomized link trajectories.
@@
Collision detection between link trajectories is a computationally expensive task, in practice we randomize each link $n_{\textrm{t}}=20$ times (and $n_{\textrm{t}}=6$ for the fruit_fly_1 network) and we rely on an efficient collision detection algorithm leveraging kd-trees [24] (See SI Sec. IV F for details).
Figure 5b shows a large variation in link confinement $C$. Across all networks, physical links are characterized by $C\approx 0$, indicating that these links are not affected by the physical proximity of other components of the network. However, we also find highly confined links with $C>10$ and even $C>100$ expected intersections, suggesting again that physicality tends to play a heterogeneous role in forming networks. In particular, the linked tree networks typically have heavy-tailed link confinement distributions (See SI Sec. V). In terms of absolute counts, the linked trees or the fruit fly neural networks show the highest values of link confinement. This can be explained by the fact that these networks are composed of multiple neurons, hence in these networks we have more complete information about the physical environment of the links, compared to networks that describe single neurons.
Figure 5: Link confinement. (a) To quantify volume exclusion at the link level, we introduce link confinement $C$, the expected number of intersections after randomizing the link trajectory. To estimate $C$ of the green link, we randomize its trajectory twice. In trial #1, the green link intersects the red link, while in in trial #2, it intersects the blue link. The blue link, however, shares and endpoint with the green link; therefore their intersection is not counted. (b) The distribution of $C$ for each physical network. There are many links with link confinement values close or equal to $C=0$, and typical values of link confinement are around $C\sim 10$, and only for the fruit fly networks, there are links with link confinement values $C>100$, indicating highly confined links for the linked tree networks.
@@ V Link confinement correlation profiles
In the previous section, we defined the link confinement $C$ as the expected number of intersections if a link would follow a random trajectory, allowing us to identify links whose trajectory is most affected by repulsive forces in the network. Here, we characterize the properties of such confined links by calculating the Kendall rank correlation $\tau$ between the link confinement $C$ and other link properties for each data set. Specifically, we focus on the (i) physical properties, complementary straightness $\bar{S}$ and link volume $V_{\textrm{link}}$ (ii) abstract network properties link betweenness $B_{\textnormal{link}}$ and link degree $k_{\textnormal{link}}(i,j)$, where the latter is defined as the sum of the degrees of the endpoints link $(i,j)$.
@@
Fig. 6 shows the correlation profiles of all 15 networks. A persistent pattern we observe across the data sets is the positive correlation between link confinement $C$ and link volume $V_{\textrm{link}}$. This is expected, since larger links have more opportunities to intersect or be intersected by neighboring links. On the other hand, correlations between link confinement and straightness show a more curious pattern: we observe that lattice-like networks tend to have positive and significant correlations $\tau$ between the link confinement $C$ and link complementary straightness $\bar{S}$, indicating that more winding links are also more confined. While for linked trees (fruit fly neural networks) we find a negative $\tau$ between $C$ and $\bar{S}$. We explain this by the fact that somata of neurons are represented as high-volume, but short and straight links in fruit fly data sets (see discussion in Sec. IV.2). This means they will be confined more due to their large volume (positive correlations between $C$ and $V$), while also following a straighter path (negative correlations between $C$ and $\bar{S}$).
@@
For correlations between link confinement $C$ and the abstract network properties, such as link betweenness $B_{\textnormal{link}}$ and degree $k_{\textnormal{link}}$, we find consistent and significant positive correlations for fruit fly neural networks. This indicates emergent correlations between the three-dimensional layout and abstract network properties of physical networks: more central links in the abstract network tend to be more confined in physical space. For lattice-like networks and trees, we find less consistent and weaker positive correlations. Overall, we are able to show that the abstract network structure and physical layout are intertwined for networks where we have sufficient information about the surrounding environment of the physical links, such as the fruit fly neural networks, which contain multiple neurons in close proximity with each other. In contrast, in cases with limited information about the we identify positive correlations for networks where
@@
Figure 6: Link confinement correlations.We compute Kendall rank correlations $\tau$ between link confinement $C$ and physical and abstract network properties of links, obtaining a correlation profile for each physical network. distinct correlation profiles for each abstract network category, which are separated into three sections separated by red lines - top for lattices, middle for trees and bottom for linked trees. Across many datasets (excluding some trees), link volume $V_{\textrm{link}}$ and link confinement $C$ has a statistically significant ($p<0.01$) small to moderate positive relationship. In addition to this general trend, there are weaker trends in terms of category-specific statistically significant correlations between link confinement $C$ and other measures - $\bar{S}$ for lattices and $B_{link}$ for trees, with a stronger trend for linked trees for all other measures - $\bar{S}$, $k_{link}$ and $B_{link}$.
@@ VI Discussion
Experimental data describing the three-dimensional shape of physical networks is increasingly becoming available, and the growth in the number and size of these data sets is expected to continue: connectome of the human brain cosists of $\approx 10^{9}$ neurons and fungal mycorrhizal networks are estimated to span $\approx 10^{17}$ km in Earth’s soil [14]. The new data calls for extending the tool set of network science to analyze, model and understand how the three-dimensional layout and physical interactions shape the structure and function of physical networks. Here, we contributed to this effort in three distinct ways: (i) We collected and standardized 15 data sets describing the three-dimensional layout of physical network from diverse domains. (ii) We characterized the structure of both the abstract network and physical layout of the 15 systems using descriptors such as the degree distribution and fractal dimension. (iii) We introduced link confinement as a method to quantify how physical interactions shape link routing in physical networks, allowing us to investigate emergent correlations between physical and abstract network properties.
@@
Our work may support future research on physical networks in several ways. First, we promote the use of labeled skeleton graphs to represent both the layout and the connectivity of physical networks. The skeleton captures the shape of the network, while the labelling identifies the physical objects corresponding the the nodes and links of the abstract network. Here, we focused on treating junction points in the skeleton as nodes and sequences of segments connecting them as links; however, the labeled skeletons are not limited to such interpretation. For example, sub-graphs representing larger functional units, such as neurons in the brain, can be identified as physical nodes.
@@
Second, our results also inform theoretical models of physical network growth. Recent work that modeled physical nodes as spheres and links between them as tubes [6, 20]. We found that most physical links follow close to straight trajectories, suggesting that linear physical network models where links are straight cylinders are indeed a useful class of models to understand physicality in networks. On the other hand, these physical network models generalize the classic Erdős-Rényi and Barabási-Albert models to physical space and thus do not restrict the node degree. We, however, found that junction points in real physical networks almost exclusively have degree three, a fact that must be accounted for by future models. Note that to obtain real physical networks with non-trivial degree distributions one must abandon identifying junction points as nodes, instead we must identify larger sub-graphs of the skeleton as physical nodes. In more formal terms, these networks can be modeled as a network-of-networks: we represent each physical node as a skeleton that has junction points with degree 3, and these physical sub-networks are bound together to form a network-of-networks with no restriction on the number of connections a sub-network can make with other sub-networks [19]. Finally, we quantified the physical confinement of individual links by comparing the path that links follow to randomized trajectories, allowing us to identify correlations between physical and abstract network properties. Future work may rely on other spatially randomized null-models to further probe the relation of the two.
@@
Our work is limited by the scope of the available data sets and computational constraints. First, our data sets do not contain information about the environment the networks are embedded in; therefore, we can only investigate interactions between the components of the network and not interactions between the networks and their surroundings. For example, we found the strongest relationship between link confinement and abstract network structure for the fruit fly neural network data sets and we found a weaker or no relationship for individual neurons. This is likely due to the fact that the fruit fly data sets contain multiple neurons, thus capture more of the environment of individual physical links. Future work may consider more complete data sets as they become available or theoretical models of network growth could incorporate non-trivial environments. Also, investigating hard-core exclusion in physical networks requires collision detection between objects with complex shapes, which is a computationally difficult problem. Future work may improve on these collision detection algorithms.

### v2

@@ IV.2 Link volume and shape
where $V_{\textnormal{seg}}(v,w)$ is the volume of each segment tracing the link $(i,j)$. We normalize the link volume by the total volume of the network $V_{\textrm{total}}$, setting the unit of measurement. We find that linked trees, or the fruit fly neural networks, have consistently high link volume heterogeneity, as their distributions span 6 to 8 orders of magnitude, which is higher compared to most lattices and trees (Fig.4a).
We also measure the aspect ratio of physical links $a(i,j)=\rho_{\textrm{link}}(i,j)/l_{\textrm{link}}(i,j)$, where $\rho_{\textrm{link}}$ is the average radius and $l_{\textrm{link}}(i,j)$ is the length of link $(i,j)$. We find the largest average aspect ratios for one of the fruit fly networks and the anthill ($\text{med}(a)\approx 0.3$), while $\text{med}(a)$ is substantially lower for other data sets (see SI S1.1). Overall, this confirms that physical links are elongated tube-like objects.
Since physical links are tube-like objects, we can capture most of their shape by characterizing their one-dimensional trajectory. Here, we calculate the deviation of the link trajectories from a straight line, quantifying how curved a link is. For this, we rely on a measure of straightness introduced originally in the context of geographical networks [31], namely we calculate the complimentary straightness for each link
@@
Calculating the median of complimentary link straightness distribution, $\text{med}(\bar{S})$, reveals that links in all of the 15 physical networks tend to follow a trajectory close to a straight line: most networks have $\text{med}(\bar{S})\approx 0.1$, meaning that the length of links is most often less than 10% longer than the optimal straight trajectory. Similarly to link volume heterogeneity, linked trees tend to cluster together and are among the networks with the straightest links with $\text{med}(\bar{S})=0.05$. Reference [26] introduced random linear physical networks, a minimal model that constructs a physical network from straight cylinders. The fact that we observed an abundance of straight or close-to-straight links lends support for using such linear physical network models to understand the role of physicality in real networks. Note, however, that although most links are close to straight, the distribution of $\bar{S}$ is right-skewed as seen in Fig. 4, which points to a smaller fraction of links that significantly deviate from a straight trajectory (see SI S3.).
@@
Finally, we computed the correlations between link straightness $S(i,j)$ and the total link length $l_{\textrm{link}}(i,j)$ and volume $V_{\textrm{link}}(i,j)$ for each data set using Kendall’s rank correlation coefficient $\tau$ [32]. Figure 4b shows that for all networks, there is a positive rank correlation $\tau>0$ between $\bar{S}(i,j)$ and $l_{\textrm{link}}(i,j)$, indicating that longer links tend to follow a more winding path. We also observe a positive correlation $\tau>0$ between $\bar{S}(i,j)$ and $V_{\textrm{link}}(i,j)$, since longer links tend to have larger volume. The only exception to this is the fruit fly neural networks, for which $\bar{S}(i,j)$ and $V_{\textrm{link}}(i,j)$ are negatively correlated $\tau<0$.
One possible cause contributing to the negative correlation is that fruit fly neural networks are composed of neurons that have large somata, which are represented in the data set as short, yet high-volume segments (see SI S1.7.).
@@
Figure 4: Link shape, length and volume. (a) Box plots of $\bar{S}$ and $V_{\textrm{link}}$, where the vertical lines indicate the median, the boxes span the $25$th-$75$th percentile range, and the whiskers extend from the $0.1$th to the $99.9$th percentile. We find that the bulk of the distributions fall in the range between $\textrm{med}(\bar{S})=0$ and $\textrm{med}(\bar{S})=0.1$, indicating that the networks are mostly composed of close-to-straight links. Link volume $V_{\textrm{link}}$ distributions span a wider range for linked trees and are the most narrow for lattices. (b) By computing Kendall rank correlation $\tau$ between link complementary straightness $\bar{S}$, link volume $V_{\textrm{link}}$ and link trajectory length $l_{\textrm{link}}$, we observe a consistent trend of $\tau>0$, meaning that longer and more voluminous links tend to have more winding paths. This trend is only reversed for the fruit fly networks, which have $\tau<0$ between $\bar{S}$ and $V_{\textrm{link}}$. This could be explained by neuron somatas, which are represented as high-volume, straight links composed of a small number of skeleton segments.
@@ IV.3 Link confinement
In the previous sections, we found that space-filling and link properties are heterogeneously distributed: most regions of space are sparse and most links are close-o-straight, yet there exists dense regions of the network and a small fraction of links follow paths that deviate from a straight line significantly. This suggests that volume exclusion or other repulsive physical interactions may also play an uneven role in shaping the network. To further investigate this hypothesis we devise a quantity that captures the confinement of a link by other components of a network.
@@
The trajectory of a physical link $(i,j)$ in our data sets is given by the ordered set ${\mathcal{T}}(i,j)$ of oriented three-dimensional segments. To generate the random trajectory ${\mathcal{T}}_{\textrm{r}}(i,j)$, we shuffle the order of the segments while maintaining their orientation and length, creating a uniform random permutation of ${\mathcal{T}}(i,j)$. The randomization preserves the endpoints and the total length of the link, but otherwise removes any correlation between subsequent segments (see SI S4.4.); therefore, the possible link trajectories ${\mathcal{T}}_{\textrm{r}}(i,j)$ have the same complimentary straightness $\bar{S}$ as the original link. Next, we estimate $I(i,j;l,k)$, the expected number of intersections between the randomized link ${\mathcal{T}}_{\textrm{r}}(i,j)$ and a non-randomized link ${\mathcal{T}}(k,l)$. To quantify the confinement of the link $(i,j)$, we sum up the expected number of intersections with other links:
@@
where the first term corresponds to intersections when link $(i,j)$ is randomized, and the second term corresponds to intersections when link $(l,k)$ is randomized. Note that the summation in Eq. (6) excludes links that share an endpoint with $(i,j)$. We do this to exclude trivial intersections from the count, since adjacent links $(i,j)$ and $(j,k)$ necessarily overlap at the junction point $j$ even for non-randomized link trajectories.
@@
Collision detection between link trajectories is a computationally expensive task, in practice we randomize each link $n_{\textrm{t}}=20$ times (and $n_{\textrm{t}}=5$ for the fruit_fly_1 network) and we rely on an efficient collision detection algorithm leveraging kd-trees [33] (see SI S4.4 and SI S4.5).
Figure 5b shows a large variation in link confinement $C$. Across all networks, physical links are characterized by $C\approx 0$, indicating that these links are not affected by the physical proximity of other components of the network. However, we also find highly confined links with $C>10$ and even $C>100$ expected intersections, suggesting again that physicality tends to play a heterogeneous role in forming networks. In particular, the linked tree networks typically have heavy-tailed link confinement distributions (see SI S5). In terms of absolute counts, the linked trees or the fruit fly neural networks show the highest values of link confinement. This can be explained by the fact that these networks are composed of multiple neurons, hence in these networks, we have more complete information about the physical environment of the links, compared to networks that describe single neurons.
Figure 5: Link confinement. (a) To quantify volume exclusion at the link level, we introduce link confinement $C$, the expected number of intersections after randomizing a link’s trajectory. To estimate $C$ of the green link, we randomize its trajectory twice. In trial #1, the green link intersects the red link, while in trial #2, it intersects the blue link. The blue link, however, shares an endpoint with the green link; therefore their intersection is not counted. (b) The distribution of $C$ for each physical network. There are many links with link confinement values close or equal to $C=0$, and typical values of link confinement are around $C\sim 10$, and only for the fruit fly networks, there are links with link confinement values $C>100$, indicating highly confined links for the linked tree networks.
@@ V Link confinement correlation profiles
In the previous section, we defined the link confinement $C$ as the expected number of intersections if a link would follow a random trajectory, allowing us to identify links whose trajectory is most affected by repulsive forces in the network. Here, we characterize the properties of such confined links by calculating the Kendall rank correlation $\tau$ between the link confinement $C$ and other link properties. Specifically, we focus on the (i) physical properties, complementary straightness $\bar{S}$ and link volume $V_{\textrm{link}}$ (ii) abstract network properties link betweenness $B_{\textnormal{link}}$ and link degree $k_{\textnormal{link}}(i,j)$, where the latter is defined as the sum of the degrees of the endpoints of link $(i,j)$.
@@
Figure 6 shows the correlation profiles of all 15 networks. A persistent pattern we observe is the positive correlation between link confinement $C$ and link volume $V_{\textrm{link}}$. This is expected since larger links have more opportunities to intersect or be intersected by neighboring links. On the other hand, correlations between link confinement and straightness show a more curious pattern: we observe that most networks tend to have positive and significant correlations $\tau$ between the link confinement $C$ and link complementary straightness $\bar{S}$, as expected, indicating that more winding links are also more confined. However, for linked trees (fruit fly neural networks) we find a negative $\tau$ between $C$ and $\bar{S}$. To explain this recall that in the fruit fly neural networks we found a negative correlation between link volume and link length and a positive correlation between straightness $\bar{S}$ (Sec. IV.2). This means that short links tend to be more confined due to their large volume, while also follow a straighter path.
@@
For correlations between link confinement $C$ and the abstract network properties, such as link betweenness $B_{\textnormal{link}}$ and degree $k_{\textnormal{link}}$, we find consistent and significant positive correlations for fruit fly neural networks. This indicates emergent correlations between the three-dimensional layout and abstract network properties of physical networks: more central links in the abstract network tend to be more confined in physical space. For lattice-like networks and trees, we find less consistent and weaker positive correlations. Overall, we are able to show that the abstract network structure and physical layout are intertwined for networks where we have sufficient information about the surrounding environment of the physical links, such as the fruit fly neural networks, which contain multiple neurons in close proximity with each other.
@@
Figure 6: Link confinement correlations. We compute the Kendall rank correlation $\tau$ between the link confinement $C$ and the physical and the abstract network properties of links, obtaining a correlation profile for each physical network. Across all datasets, except some trees, link volume $V_{\textrm{link}}$ and link confinement $C$ have a statistically significant ($p<0.01$) moderate positive relationship. We also find a significant association with complementary straightness $\bar{S}$, which is positive for lattices and trees, and negative for linked trees, consistent with the correlations between $V_{\textrm{link}}$ and $\bar{S}$ (Fig. 4). For the linked trees, which are the data sets with the most complete information about the environment of the physical links, we find a significant positive association between the centrality of the links in the abstract network (link degree $k_{\textrm{link}}$ and link betweenness $B_{\textrm{link}}$) and their link confinement $C$.
@@ VI Discussion
Experimental data describing the three-dimensional shape of physical networks is increasingly becoming available, and the growth in the number and size of these data sets is expected to continue: connectome of the human brain consists of $\approx 10^{9}$ neurons and fungal mycorrhizal networks are estimated to span $\approx 10^{17}$ km in Earth’s soil [34]. The new data calls for extending the toolset of network science to analyze, model, and understand how the three-dimensional layout and physical interactions shape the structure and function of physical networks. Here, we contributed to this effort in three distinct ways: (i) We collected and standardized 15 data sets describing the three-dimensional layout of physical networks from diverse domains. (ii) We characterized the structure of both the abstract network and physical layout of the 15 systems using descriptors such as the degree distribution and fractal dimension. (iii) We introduced link confinement as a method to quantify how physical interactions shape link trajectories in physical networks, allowing us to investigate emergent correlations between physical and abstract network properties.
@@
Our work may support future research on physical networks in several ways. First, we promote the use of labeled skeleton graphs to represent both the layout and the connectivity of physical networks. The skeleton captures the shape of the network, while the labeling identifies the physical objects corresponding the the nodes and links of the abstract network. Here, we focused on treating junction points in the skeleton as nodes and sequences of segments connecting them as links; however, the labeled skeletons are not limited to such interpretation. For example, sub-graphs representing larger functional units, such as neurons in the brain, can be identified as physical nodes.
@@
Second, our results also inform theoretical models of physical network growth. Recent work that modeled physical nodes as spheres and links between them as tubes [5, 7]. We found that most physical links follow close to straight trajectories, suggesting that linear physical network models where links are straight cylinders are indeed a useful class of models to understand physicality in networks. On the other hand, these physical network models generalize the classic Erdős-Rényi and Barabási-Albert models to physical space and thus do not restrict the node degree. We, however, found that junction points in real physical networks almost exclusively have degree three, a fact that must be accounted for by future models. Note that to obtain real physical networks with non-trivial degree distributions one must abandon identifying junction points as nodes, instead we must identify larger sub-graphs of the skeleton as physical nodes. In more formal terms, these networks can be modeled as a network-of-networks: we represent each physical node as a skeleton that has junction points with degree 3, and these physical sub-networks are bound together to form a network-of-networks with no restriction on the number of connections a sub-network can make with other sub-networks [8]. Future work may explore the relationship between the network-of-networks representation and the more fine-grained junction network representation.
Finally, we quantified the physical confinement of individual links by comparing the path that links follow to randomized trajectories, allowing us to identify correlations between physical and abstract network properties. In general, understanding the relationship between physical shape and abstract network structure is one of the key challenges of physical network research [7, 8, 35]. Future work may rely on other spatially randomized null models and abstract network measures to probe the relation between the two.
@@
Our work is limited by the scope of the available data sets and computational constraints. First, our data sets do not contain information about the environment the networks are embedded in; therefore, we can only investigate interactions between the components of the network and not interactions between the networks and their surroundings. For example, we found the strongest relationship between link confinement and abstract network structure for the fruit fly neural network data sets and we found a weaker or no relationship for individual neurons. This is likely due to the fact that the fruit fly data sets contain multiple neurons, thus capturing more of the environment of individual physical links. Future work may consider more complete data sets as they become available or theoretical models of network growth could incorporate non-trivial environments.
Data availability
All data used in the manuscript is publicly available. Code to process the data is provided at: https://github.com/lukablagoje/three-dimensional-shape-connectivity-physical-networks